leila123456:

A pizzeria makes pizzas in different sizes and with different toppings. The pizzeria wants its customers to have 12 options when they create a pizza. Which combination of size and toppings will give customers 12 options? Select three options. 4 sizes and 3 toppings 2 sizes and 10 toppings 3 sizes and 9 toppings 2 sizes and 6 toppings 3 sizes and 4 toppings

justjm:

You could try to make a sample space for each situation, and you'll eventually find that if you multiply the two then you get the total number of combinations. By that, the answers would be the ones that multiply to get a combination of 12.

justjm:

e.g., 3 sizes and 23 toppings would yield 3 * 23 = 69 options You can continue with the similar logic

leila123456:

k

leila123456:

1,4,and 5

justjm:

Yes correct bc 4*3 = 2 * 6 = 3 * 4 = 12 options

leila123456:

Thx